OpenStudy (anonymous):

calculate y'. xy^4+x^2y=x+3y

OpenStudy (anonymous):

diferentiate implicitly

OpenStudy (anonymous):

do I have to get y by itself on one side first?

OpenStudy (anonymous):

\(y^4 +xy^3y+2xy+x^2y'=1+3y'\) and now express \(y'\) as function of x and y

OpenStudy (anonymous):

I don't understand.....please help

OpenStudy (anonymous):

put one side everything that have \(y' \): \(xy^3y' +x^2y' -3y'=1-y^4-2xy\) now take \(y'\) as commun term: \(y'(xy^3+x^2-3)=1-y^4-2xy\) and just devide everything by terms in the parentesys